if a in first quadrant is=pie/19 find value of
sin(23a)-sin(3a)/sin(16a)+sin(4a)

Grade:12

2 Answers

135
10 years ago
2 sin(10a)cos(13a)/2sin(10a)cos(6a) = cos(13pie/19)/cos(6pie/19)=-1
VEERU YERRAYYA
10 years ago
=[sin (23a)-sin (3a)]/[sin (16a)+sin(4a)] =[2cos ({23a+3a}/2).sin({23a-3a}/2)] / [2sin ({16a+4a}/2).cos ({16a-4a}/2)] =cos(13a)/cos (6a) = cos(13p/19) / cos (6p/19) = -1